# Cross-Age and Cross-Site Domain Shift Impacts on Deep Learning-Based White Matter Fiber Estimation in Newborn and Baby Brains

### v1

@@ 1 Introduction
The human brain undergoes notable changes during development, particularly in white matter tracts that modulate cognitive and motor functions [1]. Accurately estimating these fibers is crucial for understanding developmental patterns and detecting abnormalities. Advances in diffusion magnetic resonance imaging (dMRI) have provided unprecedented insights into the human brain microstructure. Traditional methods, such as Constrained Spherical Deconvolution (CSD) [2] and Multi-Shell Multi-Tissue Constrained Spherical Deconvolution (MSMT-CSD) [3], have been employed to reconstruct fiber orientation distribution functions (FODs) as a proxy to the underlying microstructure. These methods often require a large number of diffusion measurements and/or multiple values, making them less feasible for uncooperative young subjects. Recently, deep learning (DL) on large datasets has allowed precise FOD estimation [4, 5, 6] with as few as six diffusion samples from developing brains [7].
@@
While DL can offer significant scanning time reduction, it is particularly faced with domain-shift problems. Such shifts can be attributed to several factors, from biological differences [8] such as age or pathologies [1] to imaging variations in protocols and scanner types (brand or field strength) [9].
@@
Data harmonization has been used for reducing variability across sites while preserving data integrity [10]. The dominant dMRI method operating at the signal level is the Rotation Invariant Spherical Harmonics (RISH) [11] that harmonizes dMRI data without model dependency but requires similar acquisition protocols and site-matched healthy controls. Deep learning techniques offer solutions to non-linear harmonization but risk overfitting and require extensive training data, potentially altering pathological information [12]. The Method of Moments (MoM) [13], which aligns diffusion-weighted imaging (DWI) features via spherical moments, stands out for its directionality preservation and independence from matched acquisition protocol or extensive training. Therefore, MoM presents a potentially beneficial approach for addressing the domain shift challenges.
@@
Furthermore, domain adaptation (DA) methods have been used to address domain shifts in medical imaging [14]. Supervised DA, particularly fine-tuning (FT) models with pre-trained weights on source domain data, is a common method, often augmented with advanced, more targeted techniques [15]. Semi-supervised DA methods, which leverage a mix of labeled and unlabeled data, can also effectively bridge domain shifts. However, both semi-supervised and unsupervised DAs face challenges in the case of significant anatomical differences, such as those between infants and neonates, where the assumption of feature space similarity may not hold.
@@
This paper investigates the domain shift effects in a DL method [7] for white matter FOD estimation in the newborn and baby populations. Our goal is to provide a detailed examination of the challenges associated with domain shifts, particularly age-related variations between these young cohorts. We propose possible solutions and emphasize the need for robust frameworks that can cope with the unique variability present in the developing brain.
@@ 2.2.1 Backbone Model
We employed the U-Net-like network [7] as the backbone for our experiments. Its proficiency lies in estimating accurate FODs from dMRI data with six diffusion directions with an extensive field of view (FoV) of , and its demonstrated accurate results when applied to dHCP newborns.
We applied the MSMT-CSD [3] using all measurements (i.e., 300 diffusion directions for dHCP and 151 directions for BCP) to generate ground-truth (GT) FODs for training and evaluation. To ensure a representative sample, subjects were randomly selected from the datasets based on the desired age range and number of subjects required by each experiment detailed in Sections 2.3 and 2.4. For each subject, we processed the diffusion signal by selecting six optimal gradient directions, normalizing, projecting onto the spherical harmonic (SH)-basis, and cropping to patches as in [7].
For each experiment, we trained the backbone network for 1000 epochs using the Adam optimizer [21] with an initial learning rate of , weight decay of , and batches of 35. We used a dropout of 0.1 to prevent overfitting. Model selection was based on the lowest mean squared error (MSE) between predicted and GT FODs in the validation set.
@@ 2.2.2 Methods for Addressing Domain Shifts
MoM [13] was employed to harmonize DWI data across sites by aligning the mean and variance using linear mapping functions . This approach adjusts each voxel’s DWI signal to the reference site’s characteristics. Median images of these moments, smoothed with a Gaussian filter to mitigate artifacts, were computed from the six optimal gradient directions and used to derive the harmonization parameters and .
@@ 2.2.3 Implementation Details and Code Availability
The training and fine-tuning were performed on an NVIDIA RTX 3090 GPU. We used TensorFlow 2.11 for our DL framework and MATLAB R2022b for MoM harmonization. The code will be made publicly available upon acceptance.
@@ 2.3 Intra-Site Age-Related Evaluation
We assessed baseline performance on the dHCP and BCP datasets, selecting 100 subjects from specified age ranges (dHCP: 29.3–44.3 post-menstrual weeks; BCP: 1.5–60 postnatal months), and allocated them into training, validation, and testing sets (70/15/15). The backbone model was trained and tested on these splits, respectively for BCP and dHCP. GT consistency was evaluated by processing two mutually exclusive subsets of the full measurements with MSMT-CSD (referred to as Gold Standards, GS), as in [7].
@@ 3.1 Intra-Site Experiments
The DL metrics are first compared to the GT consistency (GS in Table 1) for dHCP and BCP. As previously reported in [23], single-fiber predictions show good agreement, but performance decreases with multiple fibers for both datasets. This is more pronounced in three-fiber cases, which exhibit low DL performance as also reported in [7]. We therefore not consider 3-fiber metrics in subsequent experiments.
@@
Within each age group, we see similar stability in BCP compared to dHCP when training on young and testing on old subjects or vice versa. This consistency could be due to rapid development and white matter changes in the first months of life, as opposed to slower white matter development in later periods [24]. To validate this hypothesis, we explored the average white matter FA in our cohorts (Fig. 3), which indeed shows a significant increase in dHCP but a plateau in the BCP cohort. In general, AE seems to be less prone to age effect, except when training on older dHCP subjects and testing on younger ones (DL), where there is also a more pronounced decline in AR. Finally, fine-tuning on five dHCP subjects consistently reduces error rates, especially for AFD error.
@@ 5 Compliance with Ethical Standards
This research study was conducted retrospectively using human subject data made available in open access by dHCP and BCP, respectively. Ethical approval was not required, as confirmed by the licenses attached with the open-access data.
@@ 6 Acknowledgments
We acknowledge the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE, HUG, and the Leenaards and Jeantet Foundations. This research was partly funded by the Swiss National Science Foundation (grants 182602, 215641); also by the National Institute of Neurological Disorders and Stroke and the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the US National Institutes of Health (awards R01NS106030, R01NS128281, R01HD110772). We thank Hakim Ouaalam for preprocessing and parcellating BCP T2-weighted images.

### v2

@@ 2 Introduction
The human brain undergoes notable changes during development, particularly in white matter tracts that modulate cognitive and motor functions [dubois_early_2014]. Accurately estimating these fibers is crucial for understanding developmental patterns and detecting abnormalities. Advances in diffusion magnetic resonance imaging (dMRI) have provided unprecedented insights into the human brain microstructure. Traditional methods, such as Constrained Spherical Deconvolution (CSD) [tournier_robust_2007] and Multi-Shell Multi-Tissue Constrained Spherical Deconvolution (MSMT-CSD) [jeurissen_multi-tissue_2014], have been employed to reconstruct fiber orientation distribution functions (FODs) as a proxy to the underlying microstructure. These methods often require a large number of diffusion measurements and/or multiple values, making them less feasible for uncooperative young subjects. Recently, deep learning (DL) on large datasets has allowed precise FOD estimation [karimi_learning_2021, hosseini_cttrack_2022, yao_robust_2024] with as few as six diffusion samples from developing brains [kebiri_robust_2023].
@@
While DL can offer significant scanning time reduction, it is particularly faced with domain-shift problems. Such shifts can be attributed to several factors, from biological differences [bento_deep_2022] such as age or pathologies [dubois_early_2014] to imaging variations in protocols and scanner types (brand or field strength) [tax_cross-scanner_2019].
@@
Data harmonization has been used for reducing variability across sites while preserving data integrity [pinto_harmonization_2020]. The dominant dMRI method operating at the signal level is the Rotation Invariant Spherical Harmonics (RISH) [cetin_karayumak_retrospective_2019] that harmonizes dMRI data without model dependency but requires similar acquisition protocols and site-matched healthy controls. Deep learning techniques offer solutions to non-linear harmonization but risk overfitting and require extensive training data, potentially altering pathological information [bashyam_deep_2022]. The Method of Moments (MoM) [huynh_multi-site_2019], which aligns diffusion-weighted imaging (DWI) features via spherical moments, stands out for its directionality preservation and independence from matched acquisition protocol or extensive training. Therefore, MoM presents a potentially beneficial approach for addressing the domain shift challenges.
@@
Furthermore, domain adaptation (DA) methods have been used to address domain shifts in medical imaging [guan_domain_2022]. Supervised DA, particularly fine-tuning (FT) models with pre-trained weights on source domain data, is a common method, often augmented with advanced, more targeted techniques [zhou_fine-tuning_2017]. Semi-supervised DA methods, which leverage a mix of labeled and unlabeled data, can also effectively bridge domain shifts. However, both semi-supervised and unsupervised DAs face challenges in the case of significant anatomical differences, such as those between infants and neonates, where the assumption of feature space similarity may not hold.
@@
This paper investigates the domain shift effects in a DL method [kebiri_robust_2023] for white matter FOD estimation in the newborn and baby populations. Our goal is to provide a detailed examination of the challenges associated with domain shifts, particularly age-related variations between these young cohorts. We propose possible solutions and emphasize the need for robust frameworks that can cope with the unique variability present in the developing brain.
@@ 3.2.1 Backbone Model
We employed the U-Net-like network [kebiri_robust_2023] as the backbone for our experiments. Its proficiency lies in estimating accurate FODs from dMRI data with six diffusion directions with an extensive field of view (FoV) of , and its demonstrated accurate results when applied to dHCP newborns.
We applied the MSMT-CSD [jeurissen_multi-tissue_2014] using all measurements (i.e., 300 diffusion directions for dHCP and 151 directions for BCP) to generate ground-truth (GT) FODs for training and evaluation. To ensure a representative sample, subjects were randomly selected from the datasets based on the desired age range and number of subjects required by each experiment detailed in Sections 3.3 and 3.4. For each subject, we processed the diffusion signal by selecting six optimal gradient directions, normalizing, projecting onto the spherical harmonic (SH)-basis, and cropping to patches as in [kebiri_robust_2023].
For each experiment, we trained the backbone network for 1000 epochs using the Adam optimizer [kingma_adam_2015] with an initial learning rate of , weight decay of , and batches of 35. We used a dropout of 0.1 to prevent overfitting. Model selection was based on the lowest mean squared error (MSE) between predicted and GT FODs in the validation set.
@@ 3.2.2 Methods for Addressing Domain Shifts
MoM [huynh_multi-site_2019] was employed to harmonize DWI data across sites by aligning the mean and variance using linear mapping functions . This approach adjusts each voxel’s DWI signal to the reference site’s characteristics. Median images of these moments, smoothed with a Gaussian filter to mitigate artifacts, were computed from the six optimal gradient directions and used to derive the harmonization parameters and .
@@ 3.2.3 Implementation Details and Code Availability
Training and fine-tuning were performed on an NVIDIA RTX 3090 GPU. We use TensorFlow 2.11 for our DL framework and MATLAB R2022b for MoM harmonization. Our code is publicly available at http://github.com/Medical-Image-Analysis-Laboratory/dl_fiber_domain_shift.
@@ 3.3 Intra-Site Age-Related Evaluation
We assessed baseline performance on the dHCP and BCP datasets, selecting 100 subjects from specified age ranges (dHCP: 29.3–44.3 post-menstrual weeks; BCP: 1.5–60 postnatal months), and allocated them into training, validation, and testing sets (70/15/15). The backbone model was trained and tested on these splits, respectively for BCP and dHCP. GT consistency was evaluated by processing two mutually exclusive subsets of the full measurements with MSMT-CSD (referred to as Gold Standards, GS), as in [kebiri_robust_2023].
@@ 3.5 Evaluation Metrics
We quantitatively assessed FOD estimation accuracy using metrics as per [kebiri_robust_2023]: Agreement Rate (AR) for peak count consistency, Angular Error (AE) for angular discrepancy between predicted and GT FODs, and Apparent Fiber Density (AFD) from [raffelt_apparent_2012] to evaluate fiber density.
@@ 4.1 Intra-Site Experiments
The DL metrics are first compared to the GT consistency (GS in Table 1) for dHCP and BCP. As previously reported in [schilling_histological_2018], single-fiber predictions show good agreement, but performance decreases with multiple fibers for both datasets. This is more pronounced in three-fiber cases, which exhibit low DL performance as also reported in [kebiri_robust_2023]. We therefore not consider 3-fiber metrics in subsequent experiments.
@@
Within each age group, we see similar stability in BCP compared to dHCP when training on young and testing on old subjects or vice versa. This consistency could be due to rapid development and white matter changes in the first months of life, as opposed to slower white matter development in later periods [yu_differential_2020]. To validate this hypothesis, we explored the average white matter FA in our cohorts (Fig. 3), which indeed shows a significant increase in dHCP but a plateau in the BCP cohort. In general, AE seems to be less prone to age effect, except when training on older dHCP subjects and testing on younger ones (DL), where there is also a more pronounced decline in AR. Finally, fine-tuning on five dHCP subjects consistently reduces error rates, especially for AFD error.
@@ 6 Compliance with Ethical Standards
This retrospective research study used open-source human subject data from the Developing Human Connectome Project and the Baby Connectome Project, respectively, where ethical approval was not required per the data licenses.
@@ 7 Acknowledgments
We acknowledge the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE, HUG, and the Leenaards and Jeantet Foundations. This research was partly funded by the Swiss National Science Foundation (grants 182602, 215641); also by the National Institute of Neurological Disorders and Stroke and the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the US National Institutes of Health (award numbers R01NS106030, R01NS128281, R01HD110772). We thank Hakim Ouaalam for preprocessing and parcellating BCP T2-weighted images.
[title=References]